Please assist me with the equation set for this word problem.
Mary invested one amount at 8% simple interest, and a second amount at 5% interest, earning $49.16 in one year. If she had switched the amounts, she would have earned $52.37. What were the two amounts?

Answer by checkley79(3341)

.08x+.05y=49.16
.05x+.08y=52.37
Multiply the first equation by 5 & the second equation by -8 & add them.
.40x+.25y=245.80
-.40x-.64y=-418.96
-------------------
-.39y=-173.16
y=-173.16/-.39
y=$444 for one investment.
.08x+.05*444=49.16
.08x+22,20=49.16
.08x=49.16-22.20
.08x=26.96
x=26.96/.08
x=$337 for the secomd investment.
Proof:
.05*337+.08*444=52.37
16.85+35.52=52.37
52.37=52.37